The four term of a G.p is 56 while the six term is 7/8.Find the G.p and hence sum to infinity

Answer by stanbon(75887)

The four term of a G.p is 56 while the six term is 7/8.Find the G.p and hence sum to infinity
---
g(4) = a*r^3 = 56
g(6) = a*r^5 = 7/8
--------
Divide g(6) by g(4) to get r^2 = (7/8)/56 = 1/(8^2)
r = 1/8
Solve for "a"::
a*r^3 = 56
a = 56/(1/8)^3 = 28672
------
Sum to infinity:: = a/(1-r) = 28673/(7/8) = 32768
------------
Cheers,
Stan H.
-------------------